# Using Machine Learning to Explore Defect Configurations in Cd/Zn-Se/Te Compounds

### v1

@@ 7 MLFF models trained at hybrid functional accuracy
Parity plots for M3GNet-MLFF models trained on the HSE06 dataset are pictured in Figure 7(a–c), respectively for charge states $q=+1$, $q=0$, and q=1q=-1. Models for the $q=+2$ and q=2q=-2 charge states are presented in Figure S11. Each parity plot compares the HSE06-computed CFE with MLFF-predicted values across different categories: bulk (pristine supercells without defects), defects (bulk supercells containing a single defect or defect complex), defects at CdTe-ZnTe interfaces, and defects in CdTe GB structures. Despite the reduced dataset size compared to the GGA dataset, the HSE-MLFF models achieve very good accuracy with low RMSE values across different structure types. The $q=0$ test set prediction RMSE ranges from 4.8 meV/atom for bulk structures to close to 7.8 meV/atom for defect structures, 8.9 meV/atom for defects at interfaces, and 15.1 meV/atom for defects in GBs. These errors are slightly larger for $q=+1$ and q=1q=-1 defect structures but remain below 20 meV/atom for all cases, which is quite reasonable given the range of CFE values in the dataset. We also simulated selected defects in a grain boundary structure of CdTe and optimized them using the HSE-MLFF model. As illustrated in Figure S12, the MLFF achieved results comparable to full DFT but with a dramatic reduction in computational time: while conventional DFT calculations typically required around 500 minutes, the MLFF completed similar simulations in only a few minutes.
To further evaluate the MLFF’s reliability in estimating defect energetics, we compared its predictions with HSE optimization for both single defects and defect complexes (Figure S13). For neutral defect structures, the MLFF shows excellent correlation with DFT-calculated CFE, with RMSE values of 8.14 meV/atom for single defects and 9.23 meV/atom for complex defects. This suggests that the MLFF effectively captures both localized and collective defect relaxations, even for configurations with multiple defects. Figure 7(d–f) show some examples of using the HSE-MLFF model for optimizing some challenging defect configurations: the AsSe+ClSe defect complex in CdSe0.12Te0.88, a Cdi defect at the CdTe-ZnTe interface, and the AsTe substitutional defect in a CdTe GB structure. In each case, the optimization process found the ground state configuration in a few hundred steps, reaching very close to the actual HSE06 computed final energy. This gives us the confidence to use the HSE-MLFF model going forward to efficiently optimize any new defect structures at hybrid functional accuracy, before performing final HSE06+SOC calculations on the most important defects for validation and discovery. In the next section, we present some of these results, showing how high-fidelity defect formation energy diagrams could be created much faster than performing regular full DFT. Our complete workflow for predicting defect energetics using PBE/HSE-based MLFF models is illustrated in Figure 4.
@@
Figure  9: Defect formation energy diagrams for AsX, ClX, and AsX+ClX defects in (a) CdTe and (b) CdSe0.25Te0.75, under Cd-rich conditions; X = Te or Se. (c) Defect charge transition levels for AsX, ClX, and the complex AsX+ClX computed for different Se concentrations ($x=0.0$, 0.06, 0.12, 0.25) in CdSexTe1-x. Blue lines indicate the AsX (0/–1) transition level, red lines show the ClX (+1/0) transition, and purple lines show the AsX+ClX (0/–1) level. For each compound, the VBM is placed at EF = 0 eV and the CBM is placed at the band gap. All results are from HSE06+SOC calculations performed after MLFF optimization.
@@
8 HSE-MLFF Case Studies of Important Defects
@@
At this point, the “HSE-MLFF" models can be used to optimize any given defect configurations in different charge states at hybrid functional accuracy. Following this optimization, final HSE06+SOC calculations are necessary for obtaining reliable defect formation energy diagrams. Figure 8 shows the workflow of a DeFecT-FF 96 web tool we created on the nanoHUB platform to enable efficient creation of defect structures, MLFF optimization, and visualization of defect formation energy diagrams. In the next sub-sections, we present two case studies where this workflow was applied (using the HSE-MLFF models) to ultimately yield information about the relative stability and transition levels of important defects in chemical compositions of interest.
8.1 As and Cl Defects in CdSexTe1-x
Anion-site extrinsic substitutional defects were modeled in $3\times 3\times 3$ supercells of multiple CdSexTe1-x compositions (x=0, 0.06, 0.12, 0.25x=0,\,0.06,\,0.12,\,0.25). Using the Doped 86 package, we introduced AsTe, AsSe, ClTe, ClSe and the AsX+ClX double defect complexes (where $X$ denotes the preferred anion site, Te or Se). Symmetry‐breaking operations were then applied via the ShakeNBreak protocol, enabling the sampling of a diverse set of competing configurations. One example is shown in Figure S14, where 14 different symmetry-broken configurations were generated for AsSe in a CdSeTe composition. The HSE-MLFF model was then applied to identify the lowest energy configuration (configuration 5 in this case). The left panel in Figure S14 illustrates how standard DFT relaxation may lead to a false ground state, while the right panel shows the relative energies of all generated configurations, with the minimum energy structure highlighted in red.
@@ 8.1 As and Cl Defects in CdSexTe1-x
Hundreds of structures for these substitutional defects across the CdSexTe1-x compounds were relaxed with the HSE-MLFF models for different charge states until the maximum force fell below $10^{-5}$ eV/Å. Finally, single‐shot HSE06+SOC calculations were performed on the MLFF‐optimized (lowest energy) geometries to obtain accurate defect formation energy, described below for a defect D in a charge state q:
$\Delta E_{f}(D^{q})=E_{\mathrm{tot}}(D^{q})-E_{\mathrm{tot}}(\mathrm{bulk})+\sum_{i}n_{i}\,\mu_{i}+q\bigl(E_{F}+E_{\mathrm{VBM}}\bigr)+E_{\mathrm{corr}}$
@@
Here, $E_{\mathrm{tot}}(D^{q})$ and $E_{\mathrm{tot}}(\mathrm{bulk})$ are the total energies of the charged defect supercell and the pristine host supercell respectively, $n_{i}$ is the number of atoms of species $i$ removed ($n_{i}>0$) or added ($n_{i}<0$) with chemical potential $\mu_{i}$, $q$ is the defect charge state, EVBM is the computed valence band maximum energy of the bulk CdSexTe1-x compound, EF is the Fermi level which ranges across the band gap, and Ecorr is the finite‐size correction energy 6. The band gaps computed using HSE06+SOC (with a modified mixing parameter of $\alpha$=0.31) for CdTe, CdSe0.06Te0.94, CdSe0.12Te0.88, and CdSe0.25Te0.75 are respectively 1.5 eV, 1.41 eV, 138 eV, and 1.30 eV; these values are used to place the EF bounds for the defect formation energy diagrams. The charge-dependent defect formation energies additionally yield the charge transition levels as described below:
@@
This transition level marks the EF position at which charge states $q$ and $q^{\prime}$ are in equilibrium. Figure 9 presents the defect formation energy diagrams and relevant transition levels for AsX, ClX, and AsX+ClX defects across the CdSexTe1-x series, with EVBM set to 0 eV; X represents either Te or Se. Incorporation of Se is observed to deepen the AsX 0/-1 acceptor level despite the band gap going down from CdTe to CdSe0.25Te0.75, in agreement with recent experimental studies 97. The ClX +1/0 donor level remains deep in the band gap in all cases, around 1 eV from the VBM, while the AsX+ClX defect complex, interestingly, creates a 0/-1 acceptor level closer to the conduction band edge which becomes shallower with more Se content due to the lowering of the CBM. The defect energy diagrams in Figure 9(a-b) show the prevalence of the neutral state for the defect complex in the band gap, while AsX and ClX respectively create low energy acceptor and donor defects which pin the equilibrium EF around the middle of the band gap.
@@ 8.2 Native Defects and Nitrogen Impurities in ZnTe
Motivated by experimental evidence from X-ray photoelectron spectroscopy (XPS) 98, 99, 100, 101, 102, 103 indicating N incorporation in ZnTe 104, 105, 106, 107, 108, 52, 109, we employed the DeFecT-FF workflow to systematically investigate both native point defects and N-related defects in ZnTe. A $3\times 3\times 3$ ZnTe supercell was first fully relaxed using the HSE06 functional prior to defect introduction; its band gap was computed to be 2.2 eV from HSE06+SOC. The defect set included vacancies, interstitials, and antisite defects (VZn, VTe, Zni, Tei, ZnTe, TeZn), as well as different N-related defects such as Ni, NTe, and complexes Ni+Ni and NTe+Ni. To ensure thorough exploration of the potential energy landscape, we applied ShakeNBreak 110, 111 to induce perturbations, enabling the sampling of a diverse set of competing configurations. Among the hundreds of N-related configurations evaluated, the 2Ni defect emerged as the most energetically favorable, a finding further validated through additional HSE06+SOC calculations. Figure 10(a) illustrates the HSE-MLFF structural optimization and energy convergence for the 2Ni defect complex in ZnTe, and Figure 10(b) presents the HSE06+SOC computed defect formation energy diagram for N-related defects in ZnTe.
@@
Figure  10: (a) Energy as a function of optimization steps during MLFF relaxation of a ZnTe supercell with the double N interstitial (2Ni) defect complex. The inset shows the relaxed configuration with N atoms in red. (b) Defect formation energies in ZnTe under Te-rich conditions computed using HSE06+SOC on top of the HSE-MLFF optimized configurations.

### v2

@@ 3 MLFF models trained at hybrid functional accuracy
Parity plots for M3GNet-MLFF models trained on the HSE06 dataset are pictured in Figure 2(a–c), respectively for charge states $q=+1$, $q=0$, and q=1q=-1. Models for the $q=+2$ and q=2q=-2 charge states are presented in Figure S15. Each parity plot compares the HSE06-computed CFE with MLFF-predicted values across different categories: bulk (pristine supercells without defects), and defects (bulk supercells containing a single defect or defect complex). Despite the reduced dataset size compared to the GGA dataset, the HSE-MLFF models achieve very good accuracy with low RMSE values across different structure types. The $q=0$ test set prediction RMSE ranges from 4.8 meV/atom for bulk structures to 7.8 meV/atom for defect structures. These errors are similar for $q=+1$ and q=1q=-1 defect structures and remain below 12 meV/atom for all cases, which is quite reasonable given the range of CFE values in the dataset. We also simulated a limited number of CdTe dislocation core structures Fatih-CdTe-GB, Jinglong and CdTe/ZnTe interface configurations with selected defects and included them in the training dataset. The model performance for these structures is shown in Figure S14 and Figure S15. These results are not discussed in detail in the main text because the number of data points corresponding to dislocation cores and interfaces is relatively small.
Figure S16 shows the MLFF predictions for neutral defect structures separately for single defects and defect complexes, revealing low RMSE values of 8.14 meV/atom and 9.23 meV/atom respectively. This suggests that the MLFF effectively captures both localized and collective defect relaxations, even for configurations with multiple defects. Even though MLFF prediction of crystal formation energy is highly accurate for all bulk and defect structures, a true evaluation of the prediction for defect configurations involves comparing defect formation energy (and defect level) predictions from MLFF and HSE06. We used the MLFF models to optimize a selected set of defects structures. For each defect, the MLFF was used to compute the total energy entering the standard defect formation energy expression:
$\Delta E_{f}(D^{q})=E_{tot}(D^{q})-E_{tot}(bulk)+\sum_{i}n_{i}\mu_{i}+q\bigl(E_{F}+E_{VBM}\bigr)+E_{corr}$
Here, $E_{tot}(D^{q})$ and $E_{tot}(bulk)$ are the total energies of the defect and pristine supercells respectively, $n_{i}$ and $\mu_{i}$ denote the stoichiometric changes and elemental chemical potentials, $q$ is the charge state, $E_{VBM}$ is the valence band maximum energy, $E_{F}$ is the Fermi level through the band gap, and $E_{corr}$ is the correction energy Freysoldt2014-en which accounts for spurious electrostatic interactions arising from the periodic repetition of charged defects and the compensating background charge in finite supercells. In this work, $E_{corr}$ is evaluated using the Freysoldt  Freysoldt2014-en charge correction scheme, which separates long-range Coulomb interactions from short-range defect-induced potentials and aligns the electrostatic potential between defect and bulk calculations Freysoldt2014-en. Additional methodological details and convergence tests for the charge correction are provided in the SI. Importantly, the reference energy $\mu_{i}$, $E_{VBM}$ and $E_{corr}$ were evaluated from DFT and added directly to the MLFF-derived bulk and defect energies.
Figure 2(d–f) present parity plots for defect formation energy corresponding to $q=+1$ (at EF=0), $q=0$, and q=1q=-1 (at EF=0). Across the entire validation set, the RMSE in defect formation energies obtained using MLFF-optimized geometries remains below $0.20$ eV, demonstrating that the MLFFs are sufficiently accurate in capturing both structural and energetic trends for charged and neutral defects. To further assess the influence of charge corrections on MLFF-derived defect energetics, we compared three approaches: (i) using MLFF-predicted defect formation energies without any charge correction; (ii) using MLFF defect formation energies corrected using a simple average offset of $0.20$ eV for $q=+2$, $0.10$ eV for $q=+1$, $0.10$ eV for q=1q=-1, and $0.20$ eV for $q=+2$ defects; and (iii) adding known charge correction energies from DFT to MLFF-predicted defect formation energies (Figure 2(d–f)).
The comparison reveals that while uncorrected MLFF prediction shows errors close to 0.3 eV for all charge states, applying the average offset brings this error down closer to 0.2 eV which is similar to the error from adding known correction values, as listed in Table SVII. Parity plots comparing MLFF defect formation energy with DFT values for all three approaches are shown in Figure S17. Figure S18 compares defect charge transition levels predicted by the MLFF with DFT reference values. Applying an average charge correction value leads to close agreement between MLFF- and DFT-predicted defect transition levels, with RMSE values of 0.25 eV, 0.23 eV, 0.22 eV, and 0.27 eV for the $\epsilon(+2/+1)$, $\epsilon(+1/0)$, $\epsilon(0/-1)$, and $\epsilon(-1/-2)$ transitions, respectively, whereas the DFT-based charge correction values are 0.22 eV, 0.21 eV, 0.20 eV, and 0.24 eV.
To assess the reliability of our MLFF models relative to full DFT, we randomly selected 100 representative bulk and defect configurations and relaxed each structure using both methods. The DFT- and MLFF-optimized geometries were then compared using SOAP descriptors dscribe_1, dscribe_2, which provide a rotationally invariant fingerprint of the atomic environments. These high-dimensional descriptors were projected into a two-dimensional PCA PCA (principle components analysis) space, allowing direct visualization of structural similarity. In Figure S19, each DFT structure (blue) is paired with its corresponding MLFF structure (orange), with a connecting line indicating the degree of agreement. The consistently short line segments demonstrate that the MLFF reproduces DFT relaxation behavior with high accuracy. Some representative examples of MLFF-based defect structure optimizations are shown in Figure S11(d–f) and Figure S14(d–f) of the SI.
Figure  4: Benchmarking DeFecT-$F$ for a selected defect complex in CdSe0.12Te0.88. (a) Visualization of the AsSe + ClSe complex in the CdSe0.12Te0.88 alloy supercell. (b) Total energy relaxation profiles for different charge states, comparing the converged DFT energies with the DeFecT-FF–relaxed energies. (c) Defect formation energies under Cd-rich conditions for charge states +2+2 to 2-2, showing close agreement between DFT and DeFecT-FF predictions. (d) Computational cost, measured in core-hours, highlighting the significant reduction in wall-time achieved when using DeFecT-FF instead of full DFT relaxations.
@@
Figure  5: Defect formation energy diagrams for AsX, ClX, and AsX+ClX defects in (a) CdTe and (b) CdSe0.25Te0.75, under Cd-rich conditions; X = Te or Se. (c) Defect charge transition levels for AsX, ClX, and AsX+ClX, computed for different CdSexTe1-x compositions ($x=0.0$, 0.06, 0.12, 0.25). Blue lines indicate the AsX (0/–1) acceptor level, red lines show the ClX (+1/0) donor level, and purple lines show the AsX+ClX (0/–1) level. For each compound, the VBM is placed at EF = 0 eV and the CBM is placed at the value of the computed band gap. All results are from HSE06+SOC calculations performed after DeFecT-$F$ optimization.
@@
4 New Predictions with the MLFF Models: Case Studies of Important Defects
@@
The MLFF models can now be used to optimize any given defect configuration in different charge states with near–hybrid-functional accuracy. Following MLFF optimization, final HSE06+SOC single-point calculations must be performed to obtain reliable defect formation energy diagrams. Figure 3 illustrates the overall workflow of the DeFecT-FF framework defect-ff in determining the lowest energy symmetry-broken defect configuration followed by a high-fidelity understanding of the defect thermodynamics. Figure S20 shows the workflow of the DeFecT-FF defect-ff web tool we created on the nanoHUB platform to enable efficient creation of defect structures, MLFF optimization, and visualization of defect formation energy diagrams. In the next subsections, we present a few case studies demonstrating the application of this workflow to determine the relative stability and charge transition levels of important defects in selected Cd/Zn-Te/Se/S compositions which were not entirely part of the MLFF training dataset.
4.1 As+Cl Defect Complex in CdSe0.12Te0.88
As is a commonly used p-type dopant in Se-alloyed CdTe and Cl is a common impurity arising from CdCl2 treatment, which makes it important to investigate defect complexes of As and Cl in representative CdSeTe compositions. We applied the DeFecT-FF framework to simulate the AsSe + ClSe substitutional defect complex in the compound CdSe0.12Te0.88, across five possible charge states. An example configuration is illustrated in Figure 4(a); Se alloying starting from a 216-atom CdTe cubic supercell is first simulated using the special quasirandom structures (SQS) approach SQS, following which As and Cl substitution is incorporated in multiple possible locations to eventually yield the preferred combination. For each charge state between $q=+2$ and q=2q=-2, ten symmetry-broken initial structures were generated using the ShakeNBreak protocol snb_1, enabling exploration of a diverse set of competing local geometries. Each of these structures was relaxed using DeFecT-FF to identify the lowest-energy configuration prior to high-fidelity DFT refinement.
The total energy relaxation profiles for different charge states are shown in Figure 4(b). The DeFecT-FF structural optimizations converge smoothly and yield geometries that lie very close to those produced by DFT calculations. After DeFecT-FF relaxation, a single-shot HSE06+SOC calculation is performed on the predicted lowest-energy geometry for each charge state to accurately compute defect formation energies. The resulting defect formation energies (at EF=0) under Cd-rich conditions from DFT and DeFecT-FF are compared in Figure 4(c). Across all charge states, the agreement is excellent, with typical deviations well below 0.1–0.2 eV.
Thus, the DeFecT-FF geometries provide a sufficiently accurate structural foundation for defect thermodynamics for complexes in an alloyed CdSeTe composition. The computational savings are substantial: a single HSE06 relaxation of a charged defect in a $\times3$$\times$3 supercell requires approximately 512 cores multiplied by 8 to 9 hours per configuration, corresponding to a total of nearly 4096 core-hours. In contrast, the DeFecT-FF relaxations require only about $(2/60)\times 16$ core-hours per configuration, which is approximately 0.5 core-hours. The speedup therefore exceeds four orders of magnitude, as presented in Figure 4(d). We expect these trends to hold for all types of defect complexes in alloyed $\times3$$\times$3 supercells and there is confidence in DeFecT-FF reaching close agreement with hybrid DFT at a fraction of the cost.
4.2 As and Cl Defects Across CdSexTe1-x Compositions
@@
Next, we simulated multiple substitutional defects of As and Cl (including complexes) in $3\times 3\times 3$ supercells of a series of CdSexTe1-x compositions (x=0, 0.06, 0.12, 0.25x=0,\,0.06,\,0.12,\,0.25). Using the Doped doped package, we introduced defects AsTe, AsSe, ClTe, ClSe and the AsX+ClX double defect complexes (where $X$ denotes the preferred anion site, Te or Se). Symmetry‐breaking operations were then applied via the ShakeNBreak protocol, enabling the sampling of a diverse set of competing configurations. Hundreds of structures for these substitutional defects across the CdSexTe1-x compounds were relaxed with the DeFecT-FF models for different charge states until the maximum force fell below $10^{-5}$ eV/Å. Finally, single‐shot HSE06+SOC calculations were performed to obtain accurate defect formation energy.
@@
The band gaps computed using HSE06+SOC (with a modified mixing parameter of $\alpha$=0.31) for CdTe, CdSe0.06Te0.94, CdSe0.12Te0.88, and CdSe0.25Te0.75 are respectively 1.5 eV, 1.41 eV, 138 eV, and 1.30 eV; these values are used to place the EF bounds for the defect formation energy diagrams. The VBM for each composition was obtained from the bulk calculation at the HSE+SOC level using a $2\times 2\times 2$ $k$-mesh for the corresponding 216-atom $3\times 3\times 3$ supercell. The charge-dependent defect formation energies additionally yield the charge transition levels as described below:
@@
This transition level marks the EF position at which charge states $q$ and $q^{\prime}$ are in equilibrium. Figure 5 presents the defect formation energy diagrams and relevant transition levels for AsX, ClX, and AsX+ClX defects across the CdSexTe1-x series, with EVBM set to 0 eV; X represents either Te or Se. Incorporation of Se is observed to deepen the AsX 0/-1 acceptor level despite the band gap going down from CdTe to CdSe0.25Te0.75, in agreement with recent experimental studies Nardone. The ClX +1/0 donor level remains deep in the band gap in all cases, around 1 eV from the VBM, while the AsX+ClX defect complex, interestingly, creates a 0/-1 acceptor level closer to the conduction band edge which becomes shallower with more Se content due to the lowering of the CBM. The defect energy diagrams in Figure 5(a-b) show the prevalence of the neutral state for the defect complex in the band gap, while AsX and ClX respectively create low energy acceptor and donor defects which pin the equilibrium EF (obtained by applying charge-neutrality conditions) around the middle of the band gap.
@@ 4.3 Native Defects and Nitrogen Impurities in ZnTe
Motivated by experimental evidence from X-ray photoelectron spectroscopy (XPS) f__a__stevie_72130f8a, janet_mahoney_e1a9d305, artur_born_2c82645b, gustavo_lanza_7007a95e, hui_chen_dca65eb5, haipeng_xie_92ab6279 indicating N incorporation in ZnTe ji_ha_lee_5db163c4, lijuan_zhao_6c062f36, yingrui_li_bb64582f, titao_li_7535f555, daniele_dragoni_41c08f2d, eduardo_men_ndez_proupin_dae1679b, ethan_berger_ad4781d2, we employed the DeFecT-FF workflow to systematically investigate both native point defects and N-related defects in ZnTe. A $3\times 3\times 3$ ZnTe supercell (cubic zinblende phase) was first fully relaxed using the HSE06 functional prior to defect introduction; its band gap was computed to be 2.2 eV from HSE06+SOC. The defect set included vacancies, interstitials, and antisite defects (VZn, VTe, Zni, Tei, ZnTe, TeZn), as well as the following N defects: Ni, NTe, Ni+Ni, and NTe+Ni. To ensure thorough exploration of the potential energy landscape, we applied ShakeNBreak snb_1, snb_2 to induce perturbations, enabling the sampling of a diverse set of competing configurations. Among the hundreds of N-related configurations evaluated, the Ni+Ni defect complex emerged as the most energetically favorable, a finding further validated through additional HSE06+SOC calculations. Figure 6(a) illustrates the DeFecT-FF structural optimization and energy convergence for Ni+Ni in ZnTe, and Figure 6(b) presents the HSE06+SOC computed defect formation energy diagram for different N-related defects in ZnTe.
@@
Figure  6: (a) Energy as a function of optimization steps during DeFecT-$F$ relaxation of a ZnTe supercell with the double N interstitial (Ni+Ni) defect complex. The inset shows the relaxed configuration with N atoms in red. (b) Defect formation energies in ZnTe under Te-rich conditions computed using HSE06+SOC on top of the HSE-MLFF optimized configurations.
Figure  7: (a) Total energy as a function of simulation time for a microcanonical (NVE) molecular dynamics trajectory of the AsTe defect in CdTe, demonstrating excellent energy conservation over a 100 ps run. The inset shows the atomic structure before (initial) and after (final) MD simulation, with the defect site highlighted. (b) Normalized vibrational density of states (VDOS) obtained from the velocity autocorrelation function of the same trajectory, revealing dominant low-frequency vibrational modes below 4 THz that are characteristic of defect-localized and heavy-atom vibrations.
4.4 Simulating Defects at Finite Temperature
Defect formation energies in semiconductors are typically evaluated at $T=0$ K using DFT, neglecting vibrational entropy effects that can become important under realistic growth and operating conditions. While a full finite-temperature free-energy treatment is beyond the scope of this work, we demonstrate that DeFecT-FF enables finite-temperature molecular dynamics (MD) MD_1, MD_2 simulations that provide a physically grounded pathway toward incorporating such effects in future studies. As a representative example, we consider the AsTe defect in CdTe. The defect structure was first relaxed using DeFecT-FF, followed by finite-temperature microcanonical (NVE) ensemble MD simulations for the neutral charge state ($q=0$) at $T=300$ K using a 1 fs time step and a total simulation length of 100,000 steps MD_3, MD_4, MD_5. Figure 7(a) shows the energy as a function of simulation time, demonstrating excellent energy conservation over the 100 ps trajectory and confirming the numerical stability of DeFecT-FF.
Beyond validating stability, finite-temperature MD offers key physical advantages over static relaxations by enabling symmetry-breaking driven by thermal fluctuations and realistic atomic motion. This allows the system to explore the local configurational space and access dynamically stabilized or lower energy defect structures that may not be achieved at 0 K. Such MD-relaxed structures provide physically meaningful inputs for defect formation energy calculations and establish a foundation for future investigations of vibrational entropy and finite-temperature defect thermodynamics using the same framework.
During the MD simulations, atomic velocities $v_{i}(t)$ were recorded at each time step, where $i$ labels atoms in the supercell and $N$ is the total number of atoms. From these trajectories, we computed the velocity autocorrelation function as follows Rahman_Sun:
$C_{vv}(t)=\frac{1}{3N}\sum_{i=1}^{N}\langle v_{i}(0)\cdot v_{i}(t)\rangle,$ (1)
and obtained the vibrational density of states (VDOS), $g(\omega)$, using Fourier transformation:
$g(\omega)=\int_{0}^{\infty}C_{vv}(t)\,e^{-i\omega t}\,dt,$ (2)
Here, $\omega=2\pi f$ is the vibrational angular frequency. Figure 7(b) shows the normalized VDOS derived from the MD trajectory. The dominant low-frequency modes below approximately 4 THz are characteristic of defect-localized vibrations and reflect vibrational softening induced by the AsTe defect and the presence of heavy atoms in CdTe. The VDOS obtained from finite-temperature MD naturally incorporates anharmonic effects and provides direct access to vibrational entropy contributions Freysoldt2014-en which will be addressed in follow-up contributions from our group.